OpenStudy (danieke):

What is the nth term of the geometric sequence 4, 8, 16, 32, ... ? an=2(4)n−1 an=4(2)n−1 an=2(n)4 an=4(n)2 @prettygirl876

OpenStudy (danieke):

What is the common ratio of the geometric sequence whose second and fourth terms are 6 and 54, respectively? 1 2 3 4 @prettygirl876

OpenStudy (anonymous):

hello

OpenStudy (imtiaz7):

The nth term of geometric series is: an=ar^n-1 so here; a=4 and r=2 (do you know how we get it?) so equation becomes: an=4(2)^n-1

OpenStudy (imtiaz7):

@danieke so option B is your answer!

OpenStudy (danieke):

@Imtiaz7 I get it thank uu

OpenStudy (imtiaz7):

welcome :) @danieke umm you're forgetting some thing